a coil of 0.1 m * 0.1 m and of 200 turns carry current of 1 milliampere placed in uniform magnetic field of 0.1 Tesla the maximum torque that act on coil is ans. 2*10^-4

Grade:12th pass

1 Answers

Phanindra
5 years ago
Torque \tau = BiAN cos\Theta...
Theta is not given....so theta = 0...
Here, B = magnetic field = 0.1 tesla ..=10^{-1}
         i = current = 1mA = 1× 10^{-3}
          A= area = 0.1 × 0.1 = 1× 10^{-2}
        N = number of turns..= 200...
Multiply all these...
You will get the answer...Hope you have understood it...